Solve
1/3 - 5/6 = 1/x

Answer by algebrahouse.com(1659)

1/3 - 5/6 = 1/x
2/6 - 5/6 = 1/x {got common denominator}
-3/6 = 1/x {combined like terms}
-1/2 = 1/x {reduced}
x = -2 {when you have a fraction equal to 1/x, just flip both over, not off!}

For more help from me, visit: www.algebrahouse.com